A certian radioactive substance decays according to the formula A(t)=e^-0.0063t where t is time in days. What is the half life of this substance (to the nearest day)?
Answer by stanbon(48558)

A certian radioactive substance decays according to the formula
A(t) = Aoe^-0.0063t where t is time in days. What is the half life
of this substance (to the nearest day)?
--------------------------------
If you start with Ao then A(t) = (1/2)Ao will determine the halflife.
------------------
(1/2)Ao = Aoe^-0.0063t
(1/2) = e^-0.0063t
Take the natural log to get:
-0.0063t = ln(0.5)
t = ln(0.5)/-0.0063
t = 110.0233620 years
0.0233620*365 days = 8.527.. days
--------------------
Ans: Halflife = 110 yrs + 8.527 days
========================================
Cheers,
Stan H.